# Variational Inference for Sparse Poisson Regression

### v8

@@ 3.2 Continuous spike and slab prior
in which the constant $c$ is small enough to enforce sparsity to ${\boldsymbol{\beta}}_{j}$. Here, we set $c=0.001$ based on examination of various small values during cross-validation.
@@
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}=\left[S_{X}^{\boldsymbol{\xi}}+{\rm E}(\sigma^{-2})\left[\text{diag}({P^{(C)}})+c^{-1}(\text{I}-\text{diag}({P^{(C)}}))\right]\right]^{-1}$
@@
Let $D^{C}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}$. Furthermore,
@@
$\beta_{{\tau^{2}}}=\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})+\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})+{\rm E}(a^{-1}),$
and thus, ${\rm E}(\log{\tau^{2}})=\log(\beta_{{\tau^{2}}})-\psi(\alpha_{{\tau^{2}}}),$ and ${\rm E}({\tau^{-2}})=\frac{\alpha_{{\tau^{2}}}}{\beta_{{\tau^{2}}}},$
$q(a)=\text{Inv-Gamma}\left(a;1,{\rm E}({\tau^{-2}})+{A}^{-1}\right),$
and thus, ${\rm E}(\log a)=\log\left({\rm E}({\tau^{-2}})+{A}^{-1}\right)-\psi(1),$ and ${\rm E}(a^{-1})=({{\rm E}({\tau^{-2}})+{A}^{-1}})^{-1},$
@@
${P_{j}^{(C)}}=\sigma\left({\rm E}(\log\pi_{j})-{\rm E}(\log(1-\pi_{j}))-\frac{1}{2}{\rm E}({\tau^{-2}}){d_{jj}^{C}}(1-\frac{1}{c})\right),$ (12)
and ${P^{(C)}}=(1,{P_{1}^{(C)}},\ldots,{P_{p-1}^{(C)}}),$ in which $\sigma(v)=(1+\exp(-v))^{-1}$ is the sigmoid function, ${d_{jj}^{C}}=(D_{{\boldsymbol{\beta}}}^{C})_{jj}$ ,and
@@
$=$ $-M_{\boldsymbol{\xi}}^{T}(1+\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)})-\frac{1}{2}\sum_{i=1}^{n}e^{\xi_{i}}\xi_{i}^{2}-\frac{1}{2}tr(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{C})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}\allowdisplaybreaks$
$-\frac{1}{2}(1-\frac{1}{c}){\rm E}({\tau^{-2}}){\sum_{j=1}^{p-1}}{P_{j}^{(C)}}d_{jj}^{C}-\frac{1}{2c}{\rm E}({\tau^{-2}}){\sum_{j=1}^{p-1}}{d_{jj}^{C}}\allowdisplaybreaks$
$-{\left(1+\frac{p}{2}\right)}{\rm E}(\log{\tau^{2}})-\log(A^{-1}+{\rm E}({\tau^{-2}}))\allowdisplaybreaks$
$-\frac{1}{2}{\rm E}({\tau^{-2}}){\rm E}(a^{-1})+{\sum_{j=1}^{p-1}}{P_{j}^{(C)}}E\left(\frac{\log\pi_{j}}{\log(1-\pi_{j})}\right)\allowdisplaybreaks$
$+({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}(\log(1-\pi_{j}))\allowdisplaybreaks$
$-{{\rm E}(a^{-1})}/{A}+\frac{1}{2}\log|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}|-\alpha_{{\tau^{2}}}\log{\boldsymbol{\beta}}_{{\tau^{2}}}\allowdisplaybreaks$
$+\log\Gamma(\alpha_{{\tau^{2}}})+(\alpha_{{\tau^{2}}}-1){\rm E}(\log{\tau^{2}})+{\boldsymbol{\beta}}_{{\tau^{2}}}{\rm E}({\tau^{-2}})\allowdisplaybreaks$
@@
  $\bullet$ update ${\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}\leftarrow{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}),$ and ${d_{jj}^{C}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)j}^{2}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)jj}$
@@
  $\bullet$ update ${\rm E}({\tau^{-2}})\leftarrow{\dfrac{p-1}{2}}\left(\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})+\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})+{\rm E}(a^{-1})\right)^{-1},$
  $\bullet$ update ${\rm E}(a^{-1})=\left({\rm E}({\tau^{-2}})+A^{-1}\right)^{-1},$
  $\bullet$ update ${\rm E}(\log\pi_{j})=\psi({\rho}_{1}+{P_{j}^{(C)}})-\psi({\rho}_{1}+{\rho}_{2}+1)$ and ${\rm E}(\log(1-\pi_{j}))=\psi({\rho}_{2}-{P_{j}^{(C)}}+1)-\psi({\rho}_{1}+{\rho}_{2}+1)$,
  $\bullet$ update ${P_{j}^{(C)}}\leftarrow\sigma\left({\rm E}(\log\pi_{j})-{\rm E}(\log(1-\pi_{j}))-\frac{1}{2}{\rm E}({\tau^{-2}}){d_{jj}^{C}}(1-\frac{1}{c})\right),$
@@ 3.3 Bernoulli sparsity-enforcing prior
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}=\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+\text{diag}\Big{(}{\rm E}(\alpha)\Big{)}\right]^{-1},$
@@
in which $\Omega=({P^{(B)}})({P^{(B)}})^{\top}+\text{diag}({P^{(B)}})({I_{p}}-\text{diag}({P^{(B)}}))$. Let $D^{B}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}$. Furthermore, for $j=1,\ldots,{p-1}$
@@
${P_{j}^{(B)}}=\sigma\left((\mathbf{y}-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}d_{jj}^{(B)}-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\xi}d_{ij}^{B}+{\rm E}\left(\log\frac{\pi_{j}}{1-\pi_{j}}\right)\right),$ (16)
@@
$q(\alpha_{j})=\text{Gamma}\left(\alpha_{j};a_{j}+\frac{1}{2},b_{j}+\frac{d_{jj}^{B}}{2}\right),$
thus, ${\rm E}(\alpha_{j})={(a_{j}+{1}/{2})}/{(b_{j}+{d_{jj}^{B}}/{2})},$ ${\rm E}(\log\alpha_{j})=\psi(a_{j}+{1}/{2})-\log(b_{j}+{d_{jj}^{B}}/{2})$, and
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}{P^{(B)}}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}-\frac{1}{2}\text{tr}\left[D^{B}_{{\boldsymbol{\beta}}}(S_{X}^{\boldsymbol{\xi}}\odot\Omega)\right]+\allowdisplaybreaks$
$-\sum_{i=1}^{n}\big{(}e^{\xi_{i}}+\xi_{i}-\frac{\xi_{i}^{2}}{2}\big{)}+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\log\alpha_{j})\allowdisplaybreaks$
$-\frac{1}{2}\text{tr}\big{[}D^{B}_{{\boldsymbol{\beta}}}\text{diag}({\rm E}(\alpha))\big{]}+\frac{1}{2}\log\big{(}|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}|\big{)}\allowdisplaybreaks$
@@
$+{\sum_{j=1}^{p-1}}\log\Gamma({\rho}_{1}+{P_{j}^{(B)}})+{\sum_{j=1}^{p-1}}\log\Gamma({\rho}_{2}-{P_{j}^{(B)}}+1)-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\log\alpha_{j})\allowdisplaybreaks$
$-{\sum_{j=1}^{p-1}}\big{[}(a_{j}+1/2)\log(b_{j}+{d_{jj}^{B}}/2)-b_{j}\frac{a_{j}+1/2}{b_{j}+{{d_{jj}^{B}}}/{2}}\big{]}+\text{Const.}$ (17)
@@
  $\bullet$ update ${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}\leftarrow\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+\text{diag}\Big{(}{\rm E}(\alpha)\Big{)}\right]^{-1},$ $\Omega=({P^{(B)}})({P^{(B)}})^{\top}+\text{diag}({P^{(B)}})({I_{p}}-\text{diag}({P^{(B)}})),$
@@
  $\bullet$ update ${\rm E}(\alpha_{j})={(a_{j}+1/2)}/{(b_{j}+{{d_{jj}^{B}}}/{2})},$ for $j=1,\dots,{p-1},$
  $\bullet$ update ${\rm E}(\log\pi_{j})=\psi({P_{j}^{(B)}}+{\rho}_{1})-\psi({\rho}_{1}+{\rho}_{2}+1),$ and ${\rm E}(\log(1-\pi_{j}))=\psi({\rho}_{2}-{P_{j}^{(B)}}+1)-\psi({\rho}_{1}+{\rho}_{2}+1),$ for $j=1,\dots,{p-1},$
  $\bullet$ update ${P_{j}^{(B)}}=\sigma\left((\mathbf{y}-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}{d_{jj}^{B}}-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\xi}d_{ij}^{B}+{\rm E}\left(\log\frac{\pi_{j}}{1-\pi_{j}}\right)\right),$ for $j=1,\dots,{p-1},$

### v9

@@ 3.2 Continuous spike and slab prior
in which the constant $c$ is small enough to enforce sparsity to ${\boldsymbol{\beta}}_{j}$. Several sensitivity analysis was conducted to assess the influence of the hyperparameter $c$, which governs the variance of the spike component in the continuous spike-and-slab prior. The model’s predictive performance was evaluated on some real datasets (see Section 7) using the test relative prediction error for values of $c$ ranging from $0.001$ to $0.00001$. The results indicated that the predictive error was completely invariant to the choice of $c$ within this interval. This robustness suggests that the variable selection mechanism performs consistently, effectively shrinking irrelevant coefficients to zero provided $c$ is chosen sufficiently small. Based on this observed insensitivity, the value $c=0.001$ was maintained for all reported analyses.
@@
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}=\left[S_{X}^{\boldsymbol{\xi}}+{\rm E}_{q}(\sigma^{-2})\left[\text{diag}({P^{(C)}})+c^{-1}(\text{I}-\text{diag}({P^{(C)}}))\right]\right]^{-1}$
@@
Let $D^{(C)}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}$. Furthermore,
@@
$\beta_{{\tau^{2}}}=\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})+\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})+{\rm E}_{q}(a^{-1}),$
and thus, ${\rm E}_{q}(\log{\tau^{2}})=\log(\beta_{{\tau^{2}}})-\psi(\alpha_{{\tau^{2}}}),$ and ${\rm E}_{q}({\tau^{-2}})=\frac{\alpha_{{\tau^{2}}}}{\beta_{{\tau^{2}}}},$
$q(a)=\text{Inv-Gamma}\left(a;1,{\rm E}_{q}({\tau^{-2}})+{A}^{-1}\right),$
and thus, ${\rm E}_{q}(\log a)=\log\left({\rm E}_{q}({\tau^{-2}})+{A}^{-1}\right)-\psi(1),$ and ${\rm E}_{q}(a^{-1})=({{\rm E}_{q}({\tau^{-2}})+{A}^{-1}})^{-1},$
@@
${P_{j}^{(C)}}=\sigma\left({\rm E}_{q}(\log\pi_{j})-{\rm E}_{q}(\log(1-\pi_{j}))-\frac{1}{2}{\rm E}_{q}({\tau^{-2}}){d_{jj}^{(C)}}(1-\frac{1}{c})\right),$ (12)
and ${P^{(C)}}=(1,{P_{1}^{(C)}},\ldots,{P_{p-1}^{(C)}}),$ in which $\sigma(v)=(1+\exp(-v))^{-1}$ is the sigmoid function, ${d_{jj}^{(C)}}=(D_{{\boldsymbol{\beta}}}^{(C)})_{jj}$ ,and
@@
$=$ $-M_{\boldsymbol{\xi}}^{T}(1+\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)})-\frac{1}{2}\sum_{i=1}^{n}e^{\xi_{i}}\xi_{i}^{2}-\frac{1}{2}tr(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{(C)})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}\allowdisplaybreaks$
$-\frac{1}{2}(1-\frac{1}{c}){\rm E}_{q}({\tau^{-2}}){\sum_{j=1}^{p-1}}{P_{j}^{(C)}}d_{jj}^{(C)}-\frac{1}{2c}{\rm E}_{q}({\tau^{-2}}){\sum_{j=1}^{p-1}}{d_{jj}^{(C)}}\allowdisplaybreaks$
$-{\left(1+\frac{p}{2}\right)}{\rm E}_{q}(\log{\tau^{2}})-\log(A^{-1}+{\rm E}_{q}({\tau^{-2}}))\allowdisplaybreaks$
$-\frac{1}{2}{\rm E}_{q}({\tau^{-2}}){\rm E}_{q}(a^{-1})+{\sum_{j=1}^{p-1}}{P_{j}^{(C)}}E\left(\frac{\log\pi_{j}}{\log(1-\pi_{j})}\right)\allowdisplaybreaks$
$+({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log(1-\pi_{j}))\allowdisplaybreaks$
$-{{\rm E}_{q}(a^{-1})}/{A}+\frac{1}{2}\log|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}|-\alpha_{{\tau^{2}}}\log{\boldsymbol{\beta}}_{{\tau^{2}}}\allowdisplaybreaks$
$+\log\Gamma(\alpha_{{\tau^{2}}})+(\alpha_{{\tau^{2}}}-1){\rm E}_{q}(\log{\tau^{2}})+{\boldsymbol{\beta}}_{{\tau^{2}}}{\rm E}_{q}({\tau^{-2}})\allowdisplaybreaks$
@@
  $\bullet$ update ${\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}\leftarrow{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)}\mathbf{X}^{\top}(\mathbf{y}-M_{\boldsymbol{\xi}}),$ and ${d_{jj}^{(C)}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)j}^{2}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(C)jj}$
@@
  $\bullet$ update ${\rm E}_{q}({\tau^{-2}})\leftarrow{\dfrac{p-1}{2}}\left(\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})+\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})+{\rm E}_{q}(a^{-1})\right)^{-1},$
  $\bullet$ update ${\rm E}_{q}(a^{-1})=\left({\rm E}_{q}({\tau^{-2}})+A^{-1}\right)^{-1},$
  $\bullet$ update ${\rm E}_{q}(\log\pi_{j})=\psi({\rho}_{1}+{P_{j}^{(C)}})-\psi({\rho}_{1}+{\rho}_{2}+1)$ and ${\rm E}_{q}(\log(1-\pi_{j}))=\psi({\rho}_{2}-{P_{j}^{(C)}}+1)-\psi({\rho}_{1}+{\rho}_{2}+1)$,
  $\bullet$ update ${P_{j}^{(C)}}\leftarrow\sigma\left({\rm E}_{q}(\log\pi_{j})-{\rm E}_{q}(\log(1-\pi_{j}))-\frac{1}{2}{\rm E}_{q}({\tau^{-2}}){d_{jj}^{(C)}}(1-\frac{1}{c})\right),$
@@ 3.3 Bernoulli sparsity-enforcing prior
${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}=\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+\text{diag}\Big({\rm E}_{q}(\alpha)\Big)\right]^{-1},$
@@
in which $\Omega=({P^{(B)}})({P^{(B)}})^{\top}+\text{diag}({P^{(B)}})({I_{p}}-\text{diag}({P^{(B)}}))$. Let $D^{(B)}_{{\boldsymbol{\beta}}}={\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}^{\top}+{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}$. Furthermore, for $j=1,\ldots,{p-1}$
@@
${P_{j}^{(B)}}=\sigma\left((\mathbf{y}-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}d_{jj}^{(B)}-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\xi}d_{ij}^{B}+{\rm E}_{q}\left(\log\frac{\pi_{j}}{1-\pi_{j}}\right)\right),$ (16)
@@
$q(\alpha_{j})=\text{Gamma}\left(\alpha_{j};a_{j}+\frac{1}{2},b_{j}+\frac{d_{jj}^{(B)}}{2}\right),$
thus, ${\rm E}_{q}(\alpha_{j})={(a_{j}+{1}/{2})}/{(b_{j}+{d_{jj}^{(B)}}/{2})},$ ${\rm E}_{q}(\log\alpha_{j})=\psi(a_{j}+{1}/{2})-\log(b_{j}+{d_{jj}^{(B)}}/{2})$, and
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}{P^{(B)}}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}-\frac{1}{2}\text{tr}\left[D^{(B)}_{{\boldsymbol{\beta}}}(S_{X}^{\boldsymbol{\xi}}\odot\Omega)\right]+\allowdisplaybreaks$
$-\sum_{i=1}^{n}\big(e^{\xi_{i}}+\xi_{i}-\frac{\xi_{i}^{2}}{2}\big)+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\alpha_{j})\allowdisplaybreaks$
$-\frac{1}{2}\text{tr}\big[D^{(B)}_{{\boldsymbol{\beta}}}\text{diag}({\rm E}_{q}(\alpha))\big]+\frac{1}{2}\log\big(|{\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}|\big)\allowdisplaybreaks$
@@
$+{\sum_{j=1}^{p-1}}\log\Gamma({\rho}_{1}+{P_{j}^{(B)}})+{\sum_{j=1}^{p-1}}\log\Gamma({\rho}_{2}-{P_{j}^{(B)}}+1)-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\alpha_{j})\allowdisplaybreaks$
$-{\sum_{j=1}^{p-1}}\big[(a_{j}+1/2)\log(b_{j}+{d_{jj}^{(B)}}/2)-b_{j}\frac{a_{j}+1/2}{b_{j}+{{d_{jj}^{(B)}}}/{2}}\big]+\text{Const.}$ (17)
@@
  $\bullet$ update ${\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(B)}\leftarrow\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+\text{diag}\Big({\rm E}_{q}(\alpha)\Big)\right]^{-1},$ $\Omega=({P^{(B)}})({P^{(B)}})^{\top}+\text{diag}({P^{(B)}})({I_{p}}-\text{diag}({P^{(B)}})),$
@@
  $\bullet$ update ${\rm E}_{q}(\alpha_{j})={(a_{j}+1/2)}/{(b_{j}+{{d_{jj}^{(B)}}}/{2})},$ for $j=1,\dots,{p-1},$
  $\bullet$ update ${\rm E}_{q}(\log\pi_{j})=\psi({P_{j}^{(B)}}+{\rho}_{1})-\psi({\rho}_{1}+{\rho}_{2}+1),$ and ${\rm E}_{q}(\log(1-\pi_{j}))=\psi({\rho}_{2}-{P_{j}^{(B)}}+1)-\psi({\rho}_{1}+{\rho}_{2}+1),$ for $j=1,\dots,{p-1},$
  $\bullet$ update ${P_{j}^{(B)}}=\sigma\left((\mathbf{y}-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}{d_{jj}^{(B)}}-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\xi}d_{ij}^{(B)}+{\rm E}_{q}\left(\log\frac{\pi_{j}}{1-\pi_{j}}\right)\right),$ for $j=1,\dots,{p-1},$